Consider two vectors A and B. Vector A has a magnitude of 13 and points at an angle of 72 degrees above the +x axis. Vector B has a magnitude of 8 and points at an angle of 31 degrees below the +x axis.

a.     What are the x and y components of vector A?

 

 

 

 

b.     What are the x and y components of vector B?

 

 


 

 

c.     Suppose vector C is defined as C=A+B. What are the x and y components of vector C?

 

 

 

 

d.     What is the magnitude of vector C?

 

 

 

e.     At what angle does vector C point?





1
Expert's answer
2021-01-20T12:07:17-0500

Answer

a) What are the x and y component of vector A

"A_x=13cos72^\u00b0=4" unit

"A_y=13sin72^\u00b0=12.37" unit

b) What are the x and y component of vector B

"B_x=8cos31^\u00b0=6.86" unit

"B_y=8sin31^\u00b0=4.12" unit

c) C=A+B so it's components

"C_x=10.86" unit

"C_y=" 16.49unit


d) magnitude of C

"C=\sqrt{C_x^2+C_y^2}=19.72" unit

e) angle

"\theta=tan^{-1}(\frac{C_y}{C_x}) =57^°"